Solving Exponential Equation Practice:

  1. 52x = 510
  2. 53x = 1256x
  3. 2x-3 = 85
1
Expert's answer
2022-01-10T18:48:11-0500

1.

"5^{2x}=5^{10}\\2x=10\\x=5"

2.

"5^{3x}=5^{18x}\\3x=18x\\3x-18x=0\\-15x=0\\x=\frac{0}{15}\\x=0"


3.

"2^{x-3}=2^{3×5}\\x-3=15\\x=3+15\\x=18"